OpenStudy (bieber896):

what is the simplified form of each expression? (-2x^8) * 3y^9 * 2x^4 a. 3x^12y^9 b. -12x^72y^9 c. -12xy^21 d. -12x^12y^9

OpenStudy (bieber896):

@misty1212

OpenStudy (bieber896):

@imqwerty

OpenStudy (bieber896):

@Gabylovesyou

OpenStudy (ghostgate):

Well it looks like you multiply like terms over: I believe it would be like this for starters to simplify it a bit for you: (-2x^8) * 3y^9 = -6x^8y^9 ~ With that being said, what would you do with: -6x^8y^9 * 2x^4 = ?

OpenStudy (bieber896):

combine all the exponents?

OpenStudy (ghostgate):

No, multiplication with like-terms is like addition, for example: x^2 * x^2 = x^4 since it's like:(x*x)*(x*x), anyhow for this you will only be multiplying the 'base number' as I say which would be '-6' and '2' and you would be 'adding' or 'multiplying' the 'x' variables exponents which is '^8' and '^4'. Do you get what to do now?

OpenStudy (bieber896):

yeah so multiply all the base numbers first right?

OpenStudy (ghostgate):

You can, but adding the exponents at the same time isn't going to change much, your final answer is: -6x^8y^9 * 2x^4 = D. -12x^12y^9 You get that, right? I hope this helps! Have a great day! {---Ghostgate---}

OpenStudy (bieber896):

oh okay

OpenStudy (bieber896):

thanks!

OpenStudy (ghostgate):

You're welcome! ;)